The following information concerns production in the Forging Department for November. All direct materials are placed into the process at the beginning of production, and conversion costs are incurred evenly throughout the process. The beginning inventory consists of $81,780 of direct materials. ACCOUNT Work in Process—Forging Department ACCOUNT NO. Date Item Debit Credit Balance Debit Credit Nov. 1 Bal., 8,700 units, 70% completed 101,268 30 Direct materials, 79,000 units 718,900 820,168 30 Direct labor 110,200 930,368 30 Factory overhead 152,183 ? 1,082,551 30 Goods finished, ? units ? 30 Bal., 7,000 units, 70% completed ? Cost per equivalent units of $9.10 for Direct Materials and $3.30 for Conversion Costs.

Based on the above data, determine each of the following amounts. If required, round your interim calculations to two decimal places. Round final answers (a-c) to the nearest dollar.

a. Cost of beginning work in process inventory completed in November. $

b. Cost of units transferred to the next department during November. $

c. Cost of ending work in process inventory on November 30. $

d. Costs per equivalent unit of direct materials and conversion included in the November 1 beginning work in process. If required, round your answers to two decimal places. Direct materials cost per equivalent unit $ Conversion cost per equivalent unit $

e. The November increase or decrease in costs per equivalent unit for direct materials and conversion from the previous month. If required, round your answers to two decimal places. Increase or Decrease Amount Change in direct materials cost per equivalent unit Decrease $ Change in conversion cost per equivalent unit Increase
Answer #1
Equivalent units
Physical units Material Conversion
Beginning WIP 8,700 0 2,610
Units started and completed (87,700-7,000-8,700) 72,000 72,000 72,000
Units completed and transferred 80,700
Ending WIP 7,000 7,000 4,900
Total 87,700 79,000 79,510
Material Conversion
Cost added during November $718,900 $262,383
Equivalent unit 79,000 79,510
Cost per equivalent unit $9.1 $3.3

a. Cost of beginning WIP:

Beginning balance $101,268
Cost incurred during November (2,610*$3.3) 8,613
Total cost of beginning WIP $109,881

b. Cost of units transferred to next department :

Cost of beginning WIP $109,881
Cost of units started and transferred (72,000*$12.4) 892,800
Total cost of units transferred to next department $1,002,681

c. Cost of ending WIP:

Material (7,000*$9.1) $63,700
Conversion (4,900*$3.3) 16,170
Total cost of ending WIP $79,870

d. Cost per equivalent unit included in beginning WIP:

Direct material = $81,780 / 8,700 = $9.4

Conversion = $101,268-81,780 / 8,700*70%

Conversion = $19,488 / 6,090 = $3.2

e. Increase or decrease in cost per unit:

Direct material = $9.4 - 9.1 = $0.3 Decrease

Conversion = $3.2 - 3.3 = $0.1 Increase